Scenario III: Goodmark Company produces two types of birthday cards: scented and regular. During the year, 20,000 scented cards and 200,000 regular cards were produced. Goodmark has two production departments: Cutting and Printing. Direct labor hours are used to assign the overhead of Cutting and machine hours are used for Printing. The data for the two producing departments are given below:

Cutting Printing Total
Direct labor hours:
Scented 10,000 10,000 20,000
Regular 150,000 10,000 160,000
Total 160,000 20,000 180,000
Machine hours:
Scented 2,000 8,000 10,000
Regular 8,000 72,000 80,000
10,000 80,000 90,000
Overhead costs $216,000 $504,000 $720,000

Required:

1. Calculate the overhead rates for each department (round to the nearest cent):
Cutting: $ per direct labor hour
Printing: $ per machine hour
2. Calculate the overhead cost per unit (round to the nearest cent):
Scented: $
Regular: $
Answer #1
1. Calculate the overhead rates for each department (round to the nearest cent):
Cutting: $216000/160000 = 1.35 per direct labor hour
Printing: $504000/80000 = 6.30 per machine hour

2) Overhead cost per unit

Scented Regular
Cutting overhead 13500 150000*1.35 = 202500
Printing overhead 8000*6.3 = 50400 453600
Total overhead 63900 656100
Unit 20000 200000
Overhead cost per unit 3.20 3.28